OpenStudy (anonymous):

Find the points on the graph of y=secx, 0

OpenStudy (anonymous):

parallel mean they have the same slope y=4/3+2/3(x) slope of this =2/3 and the slope of the curve is the derivative of its equation find that and = with the value for the slope